# SepLLM: Accelerating Large Language Models by Compressing One Segment into One Separator

### v1

@@ 5 Concluding Remarks
In this paper, we focus on efficient neural architecture modification to address the computational and storage challenges associated with processing long contexts in large language models. From the visualization of attention maps, we find that certain separator tokens consistently contribute high attention scores. Inspired by this observation, we introduce SepLLM, a sparse attention mechanism , which focuses attention computation on initial tokens, neighboring tokens and separator tokens. To achieve real-time acceleration, we also implement hardware-efficient kernels. Our training-free studies suggest that these separator tokens effectively compress segment information, enabling efficient information retrieval. Unlike previous training-free methods, SepLLM can be incorporated into the training process, including training from scratch or post-training, thus reducing disparities between training and inference. We have conducted extensive experiments across various tasks and datasets to demonstrate SepLLM’s practical effectiveness.
@@ Appendix B The Evolution of KV Caches
To explain the dynamic design for streaming setting better, we illustrate the detailed evolution of KV caches in Figure 7. As can be seen, n and $Size_{run}$ are both periodic functions after $m_{0}$ tokens. And the average KV caches used is much less than the maximum capacity c.
@@ Appendix C Training Acceleration
We list the detailed wall-clock time per iteration and throughput in Table 9 and the speed-up ratio is 1.53.
@@
We can see that the speed of SepLLM(n=128) and SepLLM(n=64) is almost the same, which is attributed to the excellent parallelization capability of Sep-Attention module.
@@ Appendix D The Performance of Different Models
Concerning different decoder-only models, we test our SepLLM on Llama3 and Pythia backbones on PG19 test dataset (generating 64K tokens). The results are shown in Table 10 (For SepLLM, $a=4,s=64,w=256,c=800$).
@@
From above table, it can be seen that for models with similar size, setting a similar KV retention rate can yield similarly good performance.
@@
Compared to Pythia-12B, the smaller model Pythia-6.9B will have a higher perplexity if the capacity of the KV caches is the same (c=800). Therefore, it is necessary to increase c to achieve a lower perplexity close to that of Pythia-12B (c=800). On the other hand, the larger models will have lower perplexity but require a longer inference time.
@@
Falcon-40B (Almazrouei et al., 2023) is an another larger architecture we adapted for evaluation the scalability of our proposed SepLLM. The experiment results are shown in Table 12, where we set a=4,s=64,w=512/720, c=800/1024 for SepLLM, and a=4, c=800/1024 for StreamingLLM. And the conclusions are similar with previous parts.
@@
In general, whether it is the base model or the instruction-tuned model, we can condense the segment information into the corresponding Key-Value pairs of the separator tokens. To illustrate, we fine-tune Llama-3-8B-instruct and Llama-3-8B-base models (Dubey et al., 2024) on LongAlpaca dataset (Chen et al., 2024) for only 200 and 500 steps, respectively. After fine-tuning process, we take GSM8K-CoT (Cobbe et al., 2021) as the benchmark for reasoning ability evaluation. We find that both the base model and the instruction-tuned model exhibit excellent performance (matching or surpassing the vanilla model with original attention mechanism). The only difference is that for Llama-3-8B-base, we need to fine-tune for more steps to achieve such performance. In comparison, Llama-3-8B-instruct requires fewer fine-tuning steps. Even in a training-free scenario, Llama-3-8B-instruct demonstrates decent performance. This indicates that the embeddings of Llama-3-8B-instruct align with the distribution required by the SepLLM architecture better.
@@
Table 13: The comparison of SepLLM adapted to Llama-3-8B (Dubey et al., 2024) with base or instruct versions.
@@ Appendix E Needle In A Haystack
To evaluate the long-context ability of our proposed SepLLM, we take Needle In A Haystack as the benchmark and compare the performance of SepLLM and StreamingLLM. The results are shown in Figure 11,11,11,11 and SepLLM can achieve more scores compared with StreamingLLM.
@@ Appendix F Discussions on Separators
We provide the following assumptions and discussions on why keeping the KV caches corresponding to separators can maintain the performance of the original model.
@@
First, these separators (commas, periods) are extremely high-frequency tokens (both in the pre-training text and in the text generated by the language model). Thus, during the pre-training process, these separators are the most common context for all other tokens in the vocabulary. Consequently, their embeddings exhibit greater similarity, resulting in larger attention values when multiplied with other tokens. Furthermore, from a logical standpoint, it is evident that separators need to be generated by the language model very frequently. Therefore, their attention values with respect to any other token cannot be too small; otherwise, they would not be generated frequently by the language model. Tokens with larger mutual attention values will incorporate more information from other tokens. Hence, from a semantic perspective, generating a separator serves as a summarization of the current segment, naturally dividing and summarizing the contextual semantics.

### v2

@@ 5 Concluding Remarks
In this paper, we focus on efficient neural architecture modification to address the computational and storage challenges associated with processing long contexts in large language models. From the visualization of attention maps, we find that certain separator tokens consistently contribute high attention scores. Inspired by this observation, we introduce SepLLM, a sparse attention mechanism , which focuses attention computation on Initial Tokens, Neighboring Tokens and Separator Tokens. To achieve wall-clock time acceleration, we also implement hardware-efficient kernels. Our training-free studies suggest that these separator tokens effectively compress segment information, enabling efficient information retrieval. Unlike previous training-free methods, SepLLM can be incorporated into the training process, including training-from-scratch or post-training, thus reducing disparities between training and inference. We have conducted extensive experiments across various tasks and datasets to demonstrate SepLLM’s practical effectiveness.
@@ Appendix B The Evolution of KV Caches
To explain the dynamic design for streaming setting better, we illustrate the detailed evolution of KV caches in Figure 7. As can be seen, n and $Size_{run}$ are both periodic functions after $m_{0}$ tokens. And the average KV cache usage is much less than the maximum capacity c.
@@ Appendix C Training Acceleration
We list the detailed wall-clock time per iteration and throughput in Table 9. The speed-up ratio is around 1.53.
@@
We can see that the speeds of SepLLM(n=128) and SepLLM(n=64) are almost the same, which is attributed to the excellent parallelization capability of Sep-Attention module.
@@ Appendix D The Performance of Different Models
Concerning different decoder-only models, we test our SepLLM on Llama3 and Pythia backbones on PG19 test dataset (generating 64K tokens). The results are shown in Table 10 (a=4, c=800 for both SepLLM and StrmLLM. s=64,w=256 for SepLLM).
@@
From the above table, it can be seen that for models with similar size, setting a similar KV retention rate can yield similarly good performance.
@@
Compared to Pythia-12B, the smaller model Pythia-6.9B will have a higher perplexity if the entire KV cache capacity is the same (c=800). Therefore, it is necessary to increase c to achieve a lower perplexity close to that of Pythia-12B (c=800). On the other hand, the larger models will have lower perplexity but require a longer inference time.
@@
Falcon-40B (Almazrouei et al., 2023) is another larger architecture we adapted to evaluate the scalability of our proposed SepLLM. The experiment results are shown in Table 12, where we set a=4,s=64,w=512/720, c=800/1024 for SepLLM, and a=4, c=800/1024 for StreamingLLM. And the conclusions are similar to the previous parts.
@@
In general, whether it is the base model or the instruction-tuned model, we can condense the segment information into the corresponding Key-Value pairs of the separator tokens. To illustrate, we fine-tune Llama-3-8B-instruct and Llama-3-8B-base models (Dubey et al., 2024) on LongAlpaca dataset (Chen et al., 2024) for only 200 and 500 steps, respectively. After fine-tuning, we take GSM8K-CoT (Cobbe et al., 2021) as the benchmark for reasoning ability evaluation. We find that both base and instruction-tuned models exhibit excellent performance (matching or surpassing the vanilla model with original attention mechanism). The only difference is that for Llama-3-8B-base, we need to fine-tune for more steps to achieve such performance. In comparison, Llama-3-8B-instruct requires fewer fine-tuning steps. Even in a training-free scenario, Llama-3-8B-instruct demonstrates decent performance. This indicates that the embeddings of Llama-3-8B-instruct align with the distribution required by the SepLLM architecture better.
@@
Table 13: The comparison of SepLLM adapted to Llama-3-8B (Dubey et al., 2024) of base or instruct versions.
@@ Appendix E Needle In A Haystack
To evaluate the long-context information retrieval ability of our proposed SepLLM, we take Needle In A Haystack [^2] https://github.com/gkamradt/LLMTest_NeedleInAHaystack as the benchmark and compare the performance of SepLLM and StreamingLLM. The results are shown in Figure 11,11,11,11 and SepLLM can achieve more scores compared to StreamingLLM. This experiment indeed validates that SepLLM can effectively compress information of segments into the KV corresponding to separators, as even though the KV corresponding to tokens in the needle (except for possibly existing separators) are discarded, SepLLM can still retrieve the needle.
@@ Appendix F Discussions on Separators
We provide the following assumptions and discussions on why keeping the KV corresponding to separators can maintain the performance of the original model.
@@
First, these separators (commas, periods) are extremely high-frequency tokens (both in the pre-training text and in the text generated by the language model). Thus, during the pre-training process, these separators are the most common context for all other tokens in the vocabulary. Consequently, their embeddings exhibit greater similarity, resulting in larger attention values when multiplied with other tokens. Furthermore, from a logical standpoint, it is evident that separators need to be generated by the language model very frequently. Therefore, their attention values with respect to any other token cannot be too small; otherwise, they would not be generated frequently by the language model. Tokens with larger mutual attention values will incorporate more information from the other tokens. Hence, from a semantic perspective, generating a separator serves as a summarization of the current segment, naturally dividing and summarizing the contextual semantics.